SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (20th Feb, 2007)

Attendees: Al Eisner, Kim Hojeong, Selina Li, Rahmat Rahmat, Su Dong, Brandon Drummond, Arik Kreisel.

Operations

  • Paged Sunday evening due to ZPD rho plot distortion. It was during high background situation. It was normalization problem due to high background; less number of high momentum events are reconstructed. (plot)
  • Almost got paged on Saturday due to 350 damaged bins in TSF & ZPD plots. As long as it's one time thing, it's OK. (plot)
  • Hojeong is not getting email when she's paged => Selina contacted Steffen and it should be OK now.

Backgrounds

  • NTR

EMT

  • NTR

IFT

  • The IFR-LST trigger board for sextant 0 diagnozed dead. Today, spare board was in, but dead in a minute. Old board is back in. It could be a communication problem. Gigi asked to take a look at TRG plots - efficiency distribution for muon pair tracks is recommended; If it's OK, it should be flat. If there is problem, there will be dips.

GLT

  • Last repair went well. Working fine.

L3

  • There are confusion in the tracking group's L3 configuration validation. (hypernews) Arik ran with same L1 but two different L3 configurations, gave xtc files to the tracking group. They have different numbers of tracks but not sure if they are comparing same number of events either. Need more investigation in efficiency from MC, purity, resolution and fit parameters. It could be related with T0 capability.

Data Quality and Monitoring

  • Taking Run 6 data. Yields in pi0 and lambda are higher than Run 5. It's not sure if the yield is per hadronic event or per ipb. Rahmat will check this.
  • Checked photon energy distribution
  • While reprocessing Run 3 data, strip chart for TRG part is found to be empty. Some other sub-systems suffer from similar problems. It could be due to the fact that the reprocessing is running from mini. Mini doesn't have detail TRG objects in them. EMT should be there but it's not either, so maybe due to other reasons...
  • For strip chart, decided to fix the ranges. If a few points are out-of-the-range, red arrows and values will be printed. Hojeong will present the finished version next week. (not-so-perfect example, the ranges in two plots are changed in a way that there ARE some points out-of-ranges.)

Configuration

  • NTR

Collaboration Meeting talks

  • Rainer will give TRG configuration talk (plenary)
  • Debbie will give operation talk (parellel, plenary)
  • Mark will give EMT talk (parellel)
  • Arik will give L3 talk (parellel)
  • Rahmat will give DQM talk (parellel)
  • Hojeong will give "a summary on what I have been working on" talk (parellel)

This page is created by Kim Hojeong (hojeong@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.